OpenStudy (anonymous):

Use an angle sum identity to verify the identity . cos 2 0 = 2 cos^2 0 - 1

OpenStudy (anonymous):

Instead of theta, I will use x. Same proof, just easier variable to type, x. cos (2x) = cos(x+x) = cos x cos x - sinx sin x using the sum formula for cosine. so, cos (2x) = cos^2x - sin^2x. Now, since by a Pytahgoren Identity, sin^2x = 1 - cos^2x, substitute this in previous equation, so get cos(2x) = cos^2x - (1 - cos^2x) = 2 cos^2x - 1 Done.